Let μ be the true average nicotine content of a brand of cigarettes whose nicotine content may reasonably be assumed Normally distributed with σ = 0.2. Suppose the observed sample mean of the nicotine content from 32 randomly selected cigarettes: = 1.4. Find the p-value associated with the following hypothesis: Ho : μ = 1.5 vs. HA : μ 1.5 What do you conclude?
Answer #1

Given that, sample size ( n ) = 32

sample mean (ar x) = 1.4

population standard deviation (sigma) = 0.2

The null and the alternative hypotheses are,

Ho : μ = 1.5 vs. HA : μ 1,5

Test statistic is,

x-μ 1.4-1.5 Vn 7t y 32

p-value = 2 * P(Z < -2.83) = 2 * 0.0023 = 0.0046

p-value = 0.0046

Here, p-value = 0.0046 < significance level (alpha) = 0.05

So, we reject the null hypothesis ( H0).